OpenStudy (anonymous):

I need help solving this equation: -(1+7x) -6(-7-x)=36

OpenStudy (anonymous):

Multiply in first 6(-7-x) = ...

OpenStudy (anonymous):

-(1+7x) - 42- 6x = 36 -1- 7x - 42 - 6x = 32 -1 - 13x - 42 = 32 -43 - 13x = 32 - 13x = 75 -x = 5.76923 okay what did i do wrong....

OpenStudy (anonymous):

-(1+7x) +42 +6x = 36

OpenStudy (anonymous):

Your sign is wrong right at the first step of distribution !

OpenStudy (anonymous):

Any more questions?

OpenStudy (anonymous):

can you distribute it for me? I dont see what i'm doing wrong accept i go to check my answer and it doesnt work

OpenStudy (anonymous):

-1-7x +42 +6x = 36 -x = 36 - 42 + 1 = -5 => x = 5

OpenStudy (anonymous):

I still don't get it, how does that work and what's up with the => x = 5 at the end?

OpenStudy (phi):

It might be helpful to view -(1+7x) -6(-7-x)=36 as -(1+7x) + -6(-7-x)=36

OpenStudy (anonymous):

I told you that your signs is wrong at first line: -(1+7x) - 42- 6x = 36 ( wrong) Compare with: -1-7x +42 +6x = 36

OpenStudy (anonymous):

-(1+7x) + -6(-7-x) = 36 -1+ - 7x + 42 + 6x = 36 41 + - x = 36 -x = -5 x = 5 I got this answer already but when I tested it, it didn't work.

OpenStudy (phi):

5 is correct

OpenStudy (anonymous):

-(1+7(5)) + -6(-7-(5)) = 36 -(1+35) + 42 + 30 = 36 -(1 + 35) + 72 = 36 -1 + -35 + 72 = 36 -36 + 72 = 36 oh I tested it wrong. I see now, okay, thank you both so much!